# Collaborative Coded Caching for Partially Connected Networks

### v2

@@ III-A Partitioning
Figure 2: Example subnetwork where each user has the same cache profile
First, we split [^2] When we split the network, we only split the users and their nonzero links to the helpers. We don’t split the helpers. the network into $L$ different sub-networks ${\mathcal{T}}_{\ell}$ where each ${\mathcal{T}}_{\ell}$ consists only of users with cache profile $\ell$ (See Figure 2). Our aim is to partition the users in ${\mathcal{T}}_{\ell}$ into a minimum number of different sets where each user in a set can successfully decode their desired message from the joint transmission dedicated to the set. Notice that for a fully connected network, if every square submatrix $\textbf{H}^{R\times R}$ of the channel matrix of ${\mathcal{T}}_{\ell}$ is invertible as assumed in [11, 12], the minimum number of such sets as defined above is $\lceil\frac{C_{\ell}}{R}\rceil$, since the DoF is equal to the number of helpers $R$ and every $R$ number of users can be served together by a joint transmission. It is easy to see that for a partially connected sub-network ${\mathcal{T}}_{\ell}$, the minimum number of such sets will be lower bounded by $\lceil\frac{C_{\ell}}{R}\rceil$. If one considers the helpers, users with the same cache profile, and the nonzero links between them together as a bipartite graph (see Figure 2), the set of disjoint nonzero links can be seen as a matching [29]. It is well known [30] that the rank of a matrix ${\mathbf{H}}$ formed by identically 00 elements and other elements drawn independently from an i.i.d continuous distribution (e.g., i.i.d. Gaussian) is with probability $1$ equal to the maximum matching in the bipartite graph formed by row indices, and column indices, where a row index $i$ and column index $j$ have an edge $(i,j)$ if ${\mathbf{H}}_{i,j}$ is not identically zero. This implies that for any $D\leq R$ number of helpers and $D$ number of users, if there are disjoint nonzero links between $D$ helper-user pairs, for instance, between $h_{1}$ and $u_{1}$, $h_{2}$ and $u_{2}$,…, $h_{D}$ and $u_{D}$ where all $h_{d}$ and $u_{d}$ are distinct, all the $D$ users can decode their message from a joint transmission.
So, the problem is reduced to finding disjoint nonzero links between helper-user pairs. For simplicity, we will denote the partitions by only the user indexes. For instance, when we say $3-4-1-7$, this implies that there are nonzero links between $h_{1}$ and $u_{3}$, $h_{2}$ and $u_{4}$, $h_{3}$ and $u_{1}$, and $h_{4}$ and $u_{7}$ such that $u_{3}$, $u_{4}$, $u_{1}$ and $u_{7}$ can be served together. And if there is 00 in a partition, the corresponding helper isn’t assigned to a user.
@@
Consider the subnetwork in Figure 2. The channel matrix ${\mathbf{H}}$ between the $4$ helpers and users $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ can be written as ${\mathbf{H}}=\left[\begin{smallmatrix}{\mathbf{h}}_{1}[1]&0&0&0\\ {\mathbf{h}}_{2}[1]&{\mathbf{h}}_{2}[2]&0&0\\ {\mathbf{h}}_{6}[1]&{\mathbf{h}}_{6}[2]&{\mathbf{h}}_{6}[3]&0\\ 0&{\mathbf{h}}_{9}[2]&0&h_{9}[4]\end{smallmatrix}\right].$ Notice that the nonzero elements in the diagonal constitute a matching. Assume $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ request the messages $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively. Helpers can transmit ${\mathbf{x}}=[X_{1},X_{2}-X_{1}\frac{{\mathbf{h}}_{2}[1]}{{\mathbf{h}}_{2}[2]},X_{3}-X_{2}\frac{{\mathbf{h}}_{6}[2]}{{\mathbf{h}}_{6}[3]}+X_{1}\frac{{\mathbf{h}}_{2}[1]{\mathbf{h}}_{6}[2]-{\mathbf{h}}_{2}[2]{\mathbf{h}}_{6}[1]}{{\mathbf{h}}_{2}[2]{\mathbf{h}}_{6}[3]},X_{4}-X_{2}\frac{{\mathbf{h}}_{9}[2]}{{\mathbf{h}}_{9}[4]}+X_{1}\frac{{\mathbf{h}}_{2}[1]{\mathbf{h}}_{9}[2]}{{\mathbf{h}}_{2}[2]{\mathbf{h}}_{9}[4]}]^{T}$ such that ${\mathbf{H}}{\mathbf{x}}=[X_{1}{\mathbf{h}}_{1}[1],X_{2}{\mathbf{h}}_{2}[2],X_{3}{\mathbf{h}}_{6}[3],X_{4}{\mathbf{h}}_{9}[4]]^{T}$, and $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ can decode $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$, respectively. Notice that the partition $1-2-6-9$ is constructed by greedily assigning to each helper the first free user to which it has a nonzero link. It is easy to see that by continuing to assign the helpers to users greedily, one has the following partitions: $3-4-7-10$, $0-5-8-11$, and $0-0-0-12$. In this way, we have $4$ partitions. However, $\frac{C_{\ell}}{R}=\frac{12}{4}=3$, so the question is this: Can we put all the users into less than $4$ partitions?
@@
The idea is this: We first split the users in each ${\mathcal{T}}_{\ell}$ into two groups: the first group consists of users who can be served by only one helper, i.e., the users with only one nonzero link connected to a helper; the second group consists of users who can be served by more than one helper, i.e., the users with multiple nonzero links connected to distinct helpers. For instance, for the subnetwork in Figure 2, the first group consists of users $u_{1}$, $u_{4}$, $u_{5}$, $u_{7}$, $u_{8}$, $u_{11}$ and $u_{12}$, and the second group consists of $u_{2}$, $u_{3}$, $u_{6}$, $u_{9}$ and $u_{10}$. We also represent the user groups by two tables: the first group by ${\mathcal{U}}_{1}$ and the second by ${\mathcal{U}}_{2}$. In each table, the columns correspond to the helpers, and the rows correspond to the users. For ${\mathcal{U}}_{1}$, if there is a nonzero link between a user $u_{k}$ and a helper $h_{i}$, the $i$-th column of the first empty row of ${\mathcal{U}}_{1}$ is filled by the user’s index, $k$. For ${\mathcal{U}}_{2}$, if a user has nonzero links connected to multiple helpers $h_{i}$, each $i$-th column of the first empty row is filled by $k$. The first and second user groups in the subnetwork in Figure 2 are represented by Table II and Table II, respectively.
@@
Recall that our aim is to put the users into a minimum number of partitions. This means that each user must be in only one partition, and no user is left out. Every user from table ${\mathcal{U}}_{1}$ can be assigned to only one helper. However, every user from table ${\mathcal{U}}_{2}$ can be assigned to multiple helpers, and the set of helpers a user can be assigned to can be distinct. Also, the number of users in columns of table ${\mathcal{U}}_{1}$ can be different. So, depending on which helpers the users in ${\mathcal{U}}_{2}$ are assigned to may affect the number of partitions.
First, we calculate the number of users in each column of table ${\mathcal{U}}_{1}$ and put the users in each row of ${\mathcal{U}}_{1}$ into partitions. Then, we use the least cost branch and bound method to optimally assign the users in table ${\mathcal{U}}_{2}$ to helpers to be put into partitions accordingly and minimize the number of partitions. In least cost branch and bound, an objective function $f$ whose argument takes values in a finite set is minimized (or maximized), where the minimization is done without checking all the possible solutions. The algorithm consists of two main steps: Branch and Bound. In Branch, a given subset of the possible solutions is split into at least two nonempty subsets, and in Bound, a lower bound (cost) $B$ on the values $f$ takes in a subset of solutions is computed. Branch and Bound are performed iteratively to find the optimal solution. First, an upper bound $f^{max}$ on the values $f$ takes is computed heuristically, or one sets $f^{max}=\infty$. If in a Bound step $B\geq f^{max}$, then the corresponding subset of solutions is discarded. Among the different sets of subsets of possible solutions, the set with the smallest lower bound is branched first.
@@
Figure 3: Least Cost Branch and Bound Solution of the Subnetwork in Figure 2
@@ III-B Transmission
Let $G_{\ell}$ denote the number of partitions in ${\mathcal{T}}_{\ell}$. For given $g\in[G_{\ell}]$, let ${\mathbf{a}}_{\ell,g}$ denote the vector corresponding to the helpers the users are assigned to in the $g$-th partition in ${\mathcal{T}}_{\ell}$ constructed in Section III-A, let ${\mathbf{p}}_{\ell,g}$ denote the users in the partition where $U_{\ell,g}$ denotes the number of users in it. For instance, for given $g$ and $l$, if the corresponding partition is $0-2-0-1$, then ${\mathbf{a}}_{\ell,g}=[h_{2},h_{4}]$ and ${\mathbf{p}}_{\ell,g}=[u_{2},u_{1}]$. Let also $\max_{\ell}{G_{\ell}}:=G$. There will be a total of $G$ rounds, where in each round $g\in[G]$, the users in partitions ${\mathbf{p}}_{\ell,g}$ for each $\ell\in[L]$ are served if the corresponding partition is not empty. For given ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$, let ${\mathcal{T}}^{g}$ denote the set ${\mathcal{T}}$ excluding the cache profiles $\ell$ where ${\mathbf{p}}_{\ell,g}$ is empty in round $g$. Then in round $g$, for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$ where ${\mathcal{T}}^{g}$ is nonempty [^3] Notice that there are $\binom{L}{\gamma L+1}-\binom{e(g)}{\gamma L+1}$ number of ${\mathcal{T}}^{g}$ that is nonempty where $e(g)$ is the number of empty partitions ${\mathbf{p}}_{\ell,g}$, for $l\in[L]$. , the following vector is transmitted
${\mathbf{x}}({\mathcal{T}})=\sum\nolimits_{\ell\in{\mathcal{T}}^{g}}z({\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}{\mathbf{w}}_{({\mathbf{p}}_{\ell,g})})$ (4)
where ${\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}$ denotes the inverse of the channel matrix between the helpers in ${\mathbf{a}}_{\ell,g}$ and the users in ${\mathbf{p}}_{\ell,g}$, ${\mathbf{w}}_{({\mathbf{p}}_{\ell,g})}$ denotes the column vector where its $k$th element is $\𝒯W^{d_{{\mathbf{p}}_{\ell,g}[k]}}_{{\mathcal{T}}\backslash\{\ell\}}$ for $k\in[U_{\ell,g}]$, and $z$ denotes the zero-padding function that adds zeros in the helper indexes not in ${\mathbf{a}}_{\ell,g}$ to the vector in its argument. For instance, if ${\mathbf{a}}_{\ell,g}=[h_{2},h_{4}]$, then ${\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}{\mathbf{w}}_{({\mathbf{p}}_{\ell,g})}$ is a $2\times 1$ vector, say $[V_{2},V_{4}]^{T}$. Then $z([V_{2},V_{4}]^{T})=[0,V_{2},0,V_{4}]^{T}$. Decoding follows similarly as the decoding for the transmission in (2).
@@
Consider the subnetwork in Figure 2, and without loss of generality, assume that the subnetwork corresponds to users with cache profile $1$, $L=3$, $t=1$ and $K=22$. For simplicity, we only write the indexes of the users and helpers in the partition vectors, and if a partition is full, we suppress the helper vector in the notation. For the second round, we have ${\mathbf{p}}_{1,2}=[2,5,8,12]$, assume also ${\mathbf{p}}_{2,2}=[14,18]$ where ${\mathbf{a}}_{2,2}=[2,4]$ and ${\mathbf{p}}_{3,2}$ is empty. There are $3$ subsets of $[L]=[3]$ of size $t+1=2$: ${\mathcal{T}}_{1}=\{1,2\}$, ${\mathcal{T}}_{2}=\{1,3\}$, and ${\mathcal{T}}_{3}=\{2,3\}$. Then, in the second round, ${\mathcal{T}}_{1}^{2}=\{1,2\}$, ${\mathcal{T}}_{2}^{2}=\{1\}$ and ${\mathcal{T}}_{3}^{2}=\{2\}$ such that ${\mathbf{x}}(\{1,2\})={\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{2\}},W^{d_{u_{5}}}_{\{2\}},W^{d_{u_{8}}}_{\{2\}},W^{d_{u_{12}}}_{\{2\}}]^{T}+z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{1\}},W^{d_{u_{18}}}_{\{1\}}]^{T})$, ${\mathbf{x}}(\{1,3\})={\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{3\}},W^{d_{u_{5}}}_{\{3\}},W^{d_{u_{8}}}_{\{3\}},W^{d_{u_{12}}}_{\{3\}}]^{T}$, ${\mathbf{x}}(\{2,3\})=z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{3\}},W^{d_{u_{18}}}_{\{3\}}]^{T})$. User $u_{14}$ can remove ${\mathbf{h}}_{14}^{T}{\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{2\}},W^{d_{u_{5}}}_{\{2\}},W^{d_{u_{8}}}_{\{2\}},W^{d_{u_{12}}}_{\{2\}}]^{T}$ since it has cache profile $2$, and decode its desired subfile $W^{d_{u_{14}}}_{\{1\}}$, as ${\mathbf{h}}_{14}^{T}z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{1\}},W^{d_{u_{18}}}_{\{1\}}]^{T})=W^{d_{u_{14}}}_{\{1\}}$. Similarly, ${\mathbf{h}}_{14}^{T}z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{3\}},W^{d_{u_{18}}}_{\{3\}}]^{T})=W^{d_{u_{14}}}_{\{3\}}$ such that $u_{14}$ can decode its requested file $W^{d_{u_{14}}}$. In a similar way, all the other users can decode their requested files.

### v3

@@ III-A Partitioning
Figure 2: Example subnetwork where each user has the same cache profile.
First, we split [^2] When we split the network, we only split the users and their nonzero links to the helpers. We don’t split the helpers. the network into $L$ different sub-networks ${\mathcal{T}}_{\ell}$. Each ${\mathcal{T}}_{\ell}$ consists of $C_{\ell}$ users, each having cache profile $\ell$ (see Figure 2). Our aim is to partition the users in ${\mathcal{T}}_{\ell}$ into the minimum number of different sets, where each user in a set can successfully decode its desired message from the joint transmission dedicated to the set. Notice that for a fully connected network, if every square submatrix $\textbf{H}^{E\times E}$ of the channel matrix of ${\mathcal{T}}_{\ell}$ is invertible as assumed in [11, 12], the minimum number of such sets as defined above is $\lceil\frac{C_{\ell}}{E}\rceil$, since the sum-DoF is equal to the number of helpers $E$ and every $E$ number of users can be served together by a joint transmission. It is easy to see that for a partially connected sub-network ${\mathcal{T}}_{\ell}$, the minimum number of such sets will be lower bounded by $\lceil\frac{C_{\ell}}{E}\rceil$. If one considers the helpers, users with the same cache profile, and the nonzero links between them together as a bipartite graph (see Figure 2), the set of disjoint nonzero links can be seen as a matching [29]. It is well known [30] that the rank of a matrix ${\mathbf{H}}$ formed by identically 00 elements and other elements drawn independently from an i.i.d continuous distribution (e.g., i.i.d. Gaussian) is with probability $1$ equal to the maximum matching in the bipartite graph formed by row indices, and column indices, where a row index $i$ and column index $j$ have an edge $(i,j)$ if ${\mathbf{H}}_{i,j}$ is not identically zero. This implies that for any $D\leq E$ number of helpers and $D$ number of users, if there are disjoint nonzero links between $D$ helper-user pairs, for instance, between $e_{1}$ and $u_{1}$, $e_{2}$ and $u_{2}$,…, $e_{D}$ and $u_{D}$ where all $e_{d}$ and $u_{d}$ are distinct, all the $D$ users can decode their message from a joint transmission.
So, the problem is reduced to finding disjoint nonzero links between helper-user pairs. For simplicity, we will denote the partitions by only the user indexes. For instance, when we say $3-4-1-7$, this implies that there are nonzero links between $e_{1}$ and $u_{3}$, $e_{2}$ and $u_{4}$, $e_{3}$ and $u_{1}$, and $e_{4}$ and $u_{7}$ such that $u_{3}$, $u_{4}$, $u_{1}$ and $u_{7}$ can be served together. And if there is 00 in a partition, the corresponding helper isn’t assigned to a user.
@@
Consider the subnetwork in Figure 2. The channel matrix ${\mathbf{H}}$ between the $4$ helpers and users $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ can be written as ${\mathbf{H}}=\left[\begin{smallmatrix}{\mathbf{h}}_{1}[1]&0&0&0\\ {\mathbf{h}}_{2}[1]&{\mathbf{h}}_{2}[2]&0&0\\ {\mathbf{h}}_{6}[1]&{\mathbf{h}}_{6}[2]&{\mathbf{h}}_{6}[3]&0\\ 0&{\mathbf{h}}_{9}[2]&0&h_{9}[4]\end{smallmatrix}\right].$ Notice that the nonzero elements in the diagonal constitute a matching. Assume $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ request the messages $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively. Helpers can transmit ${\mathbf{x}}=[X_{1},X_{2}-X_{1}\frac{{\mathbf{h}}_{2}[1]}{{\mathbf{h}}_{2}[2]},X_{3}-X_{2}\frac{{\mathbf{h}}_{6}[2]}{{\mathbf{h}}_{6}[3]}+X_{1}\frac{{\mathbf{h}}_{2}[1]{\mathbf{h}}_{6}[2]-{\mathbf{h}}_{2}[2]{\mathbf{h}}_{6}[1]}{{\mathbf{h}}_{2}[2]{\mathbf{h}}_{6}[3]},X_{4}-X_{2}\frac{{\mathbf{h}}_{9}[2]}{{\mathbf{h}}_{9}[4]}+X_{1}\frac{{\mathbf{h}}_{2}[1]{\mathbf{h}}_{9}[2]}{{\mathbf{h}}_{2}[2]{\mathbf{h}}_{9}[4]}]^{T}$ such that ${\mathbf{H}}{\mathbf{x}}=[X_{1}{\mathbf{h}}_{1}[1],X_{2}{\mathbf{h}}_{2}[2],X_{3}{\mathbf{h}}_{6}[3],X_{4}{\mathbf{h}}_{9}[4]]^{T}$, and $u_{1}$, $u_{2}$, $u_{6}$ and $u_{9}$ can decode $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$, respectively. Notice that the partition $1-2-6-9$ is constructed by greedily assigning to each helper the first free user to which it has a nonzero link. It is easy to see that by continuing to assign the helpers to users greedily, one has the following partitions: $3-4-7-10$, $0-5-8-11$, and $0-0-0-12$. In this way, we have $4$ partitions. However, $\frac{C_{\ell}}{E}=\frac{12}{4}=3$, so the question is: Can we put all the users into less than $4$ partitions?
@@
The idea is this: We first split the users in each ${\mathcal{T}}_{\ell}$ into two groups: the first group consists of users who can be served by only one helper, i.e., the users with only one nonzero link connected to a helper; the second group consists of users who can be served by more than one helper, i.e., the users with multiple nonzero links connected to distinct helpers. For instance, for the subnetwork in Figure 2, the first group consists of users $u_{1}$, $u_{4}$, $u_{5}$, $u_{7}$, $u_{8}$, $u_{11}$ and $u_{12}$, and the second group consists of $u_{2}$, $u_{3}$, $u_{6}$, $u_{9}$ and $u_{10}$. We also represent the user groups by two tables: the first group by ${\mathcal{U}}_{1}$ and the second by ${\mathcal{U}}_{2}$. In each table, the columns correspond to the helpers, and the rows correspond to the users. For ${\mathcal{U}}_{1}$, if there is a nonzero link between a user $u_{k}$ and a helper $e_{i}$, the $i$-th column of the first empty row of ${\mathcal{U}}_{1}$ is filled by the user’s index, $k$. For ${\mathcal{U}}_{2}$, if a user has nonzero links connected to multiple helpers $e_{i}$, each $i$-th column of the first empty row is filled by $k$. The first and second user groups in the subnetwork in Figure 2 are represented by Table II and Table II, respectively.
@@
Recall that our aim is to put users into the minimum number of partitions. This means that each user must be in only one partition, and no user is left out. Every user from table ${\mathcal{U}}_{1}$ can be assigned to only one helper. However, every user from table ${\mathcal{U}}_{2}$ can be assigned to multiple helpers, and the set of helpers a user can be assigned to can be distinct. Also, the number of users in columns of table ${\mathcal{U}}_{1}$ can be different. So, depending on which helpers the users in ${\mathcal{U}}_{2}$ are assigned to may affect the number of partitions.
First, we calculate the number of users in each column of table ${\mathcal{U}}_{1}$ and put the users in each row of ${\mathcal{U}}_{1}$ into partitions. Then, we use the least cost branch and bound method to assign the users in table ${\mathcal{U}}_{2}$ to helpers to be put into partitions accordingly and minimize the number of partitions. In least cost branch and bound, an objective function $f$ whose argument takes values in a finite set is minimized (or maximized), where the minimization is done without checking all the possible solutions.
The algorithm consists of two main steps: Branch and Bound. In Branch, a given subset of the possible solutions is split into at least two nonempty subsets, and in Bound, a lower bound (cost) $B$ on the values $f$ takes in a subset of solutions is computed. Branch and Bound are performed iteratively to find the optimal solution. First, an upper bound $f^{max}$ on the values $f$ takes is computed heuristically, or one sets $f^{max}=\infty$. If in a Bound step $B\geq f^{max}$, then the corresponding subset of solutions is discarded. Among the different sets of subsets of possible solutions, the set with the smallest lower bound is branched first.
@@
Figure 3: Least cost branch and bound solution of the subnetwork in Figure 2. Each vertex corresponds to assigning a user to a helper, while an edge corresponds to the number of partitions after such an assignment.
@@ III-B Transmission
Let $G_{\ell}$ denote the number of partitions in ${\mathcal{T}}_{\ell}$. For given $g\in[G_{\ell}]$, let ${\mathbf{a}}_{\ell,g}$ denote the vector corresponding to the helpers the users are assigned to in the $g$-th partition in ${\mathcal{T}}_{\ell}$ constructed in Section III-A, let ${\mathbf{p}}_{\ell,g}$ denote the users in the partition where $U_{\ell,g}$ denotes the number of users in it. For instance, for given $g$ and $l$, if the corresponding partition is $0-2-0-1$, then ${\mathbf{a}}_{\ell,g}=[e_{2},e_{4}]$ and ${\mathbf{p}}_{\ell,g}=[u_{2},u_{1}]$. Let also $\max_{\ell}{G_{\ell}}:=G$. There will be a total of $G$ rounds, where in each round $g\in[G]$, the users in partitions ${\mathbf{p}}_{\ell,g}$ for each $\ell\in[L]$ are served if the corresponding partition is not empty. For given ${\mathcal{T}}_{j}\subseteq[L]$ of size $|{\mathcal{T}}_{j}|=\gamma L+1$, $j\in[\binom{L}{\gamma L+1}]$, let ${\mathcal{T}}^{g}_{j}$ denote the set ${\mathcal{T}}_{j}$ excluding the cache profiles $\ell$ where ${\mathbf{p}}_{\ell,g}$ is empty in round $g$. Then in round $g$, for each such subset ${\mathcal{T}}_{j}\subseteq[L]$ where ${\mathcal{T}}^{g}_{j}$ is nonempty [^3] Notice that there are $\binom{L}{\gamma L+1}-\binom{v(g)}{\gamma L+1}$ number of ${\mathcal{T}}^{g}$ that is nonempty where $v(g)$ is the number of empty partitions ${\mathbf{p}}_{\ell,g}$, for $l\in[L]$. , the following vector is transmitted
${\mathbf{x}}({\mathcal{T}}_{j})=\sum\nolimits_{\ell\in{\mathcal{T}}_{j}^{g}}z({\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}{\mathbf{w}}_{({\mathbf{p}}_{\ell,g})})$ (4)
where ${\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}$ denotes the inverse of the channel matrix between the helpers in ${\mathbf{a}}_{\ell,g}$ and the users in ${\mathbf{p}}_{\ell,g}$, ${\mathbf{w}}_{({\mathbf{p}}_{\ell,g})}$ denotes the column vector where its $k$th element is $W^{d_{{\mathbf{p}}_{\ell,g}[k]}}_{{\mathcal{T}}_{j}\backslash\{\ell\}}$ for $k\in[U_{\ell,g}]$, and $z$ denotes the zero-padding function that adds zeros in the helper indexes not in ${\mathbf{a}}_{\ell,g}$ to the vector in its argument. For instance, if ${\mathbf{a}}_{\ell,g}=[e_{2},e_{4}]$, then ${\mathbf{H}}_{({\mathbf{a}}_{\ell,g},{\mathbf{p}}_{\ell,g})}^{-1}{\mathbf{w}}_{({\mathbf{p}}_{\ell,g})}$ is a $2\times 1$ vector, say $[V_{2},V_{4}]^{T}$. Then $z([V_{2},V_{4}]^{T})=[0,V_{2},0,V_{4}]^{T}$.
@@
Consider the subnetwork in Figure 2, and without loss of generality, assume that the subnetwork corresponds to users with cache profile $1$, $L=3$, $t=1$. For simplicity, we only write the indices of the users and helpers in the partition vectors, and if a partition is full, we suppress the helper vector in the notation. For the second round, we have ${\mathbf{p}}_{1,2}=[2,5,8,12]$, assume also ${\mathbf{p}}_{2,2}=[14,18]$ where ${\mathbf{a}}_{2,2}=[2,4]$ and ${\mathbf{p}}_{3,2}$ is empty. There are $3$ subsets of $[L]=[3]$ of size $t+1=2$: ${\mathcal{T}}_{1}=\{1,2\}$, ${\mathcal{T}}_{2}=\{1,3\}$, and ${\mathcal{T}}_{3}=\{2,3\}$. Then, in the second round, ${\mathcal{T}}_{1}^{2}=\{1,2\}$, ${\mathcal{T}}_{2}^{2}=\{1\}$ and ${\mathcal{T}}_{3}^{2}=\{2\}$ such that ${\mathbf{x}}(\{1,2\})={\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{2\}},W^{d_{u_{5}}}_{\{2\}},W^{d_{u_{8}}}_{\{2\}},W^{d_{u_{12}}}_{\{2\}}]^{T}+z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{1\}},W^{d_{u_{18}}}_{\{1\}}]^{T})$, ${\mathbf{x}}(\{1,3\})={\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{3\}},W^{d_{u_{5}}}_{\{3\}},W^{d_{u_{8}}}_{\{3\}},W^{d_{u_{12}}}_{\{3\}}]^{T}$, ${\mathbf{x}}(\{2,3\})=z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{3\}},W^{d_{u_{18}}}_{\{3\}}]^{T})$. User $u_{14}$ can remove ${\mathbf{h}}_{14}^{T}{\mathbf{H}}^{-1}_{([2,5,8,12])}[W^{d_{u_{2}}}_{\{2\}},W^{d_{u_{5}}}_{\{2\}},W^{d_{u_{8}}}_{\{2\}},W^{d_{u_{12}}}_{\{2\}}]^{T}$ since it has cache profile $2$, and decode its desired subfile $W^{d_{u_{14}}}_{\{1\}}$, as ${\mathbf{h}}_{14}^{T}z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{1\}},W^{d_{u_{18}}}_{\{1\}}]^{T})=W^{d_{u_{14}}}_{\{1\}}$. Similarly, ${\mathbf{h}}_{14}^{T}z({\mathbf{H}}^{-1}_{([2,4],[14,18])}[W^{d_{u_{14}}}_{\{3\}},W^{d_{u_{18}}}_{\{3\}}]^{T})=W^{d_{u_{14}}}_{\{3\}}$ such that $u_{14}$ can decode its requested file $W^{d_{u_{14}}}$. In a similar way, all the other users can decode their requested files.